OpenStudy (anonymous):

an = cos(npi/2). If convergent find limit of sequence?

OpenStudy (anonymous):

an n goes to infinity?

OpenStudy (anonymous):

as*

OpenStudy (anonymous):

Determine whether the sequence is convergent or divergent. If it is convergent, find its limit. (a) an = cos(npi/2) This was the question.

OpenStudy (anonymous):

so the sequence is: \[\cos(\frac{\pi}{2}), \cos(\pi), \cos(\frac{3\pi}{2}), \cos(2\pi), \cos(\frac{5\pi}{2}), \ldots\] Do konw the value of those cos expressions?

OpenStudy (anonymous):

ugh >.< do you know* the value of those cos expressions.

OpenStudy (anonymous):

No dude. I have a midterm tomorrow and I have no Idea on how do solve this type of questions. I posted the exact question that was in the review sheet. It is tested on sequences from calc 2.

OpenStudy (anonymous):

Will you have access to a calculator during the test? its something you can type right in. If not, you will have to have your trig values memorized.

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

you dont need to know what the values are, only understand how the cos function behaves

OpenStudy (anonymous):

Your help is much appreciated. We can use a scientific calculator but not a graphing one.

OpenStudy (anonymous):

well, I take that back. in second term calc, you really should know what those values are

OpenStudy (anonymous):

1,0,-1,0,1,0,-1, etc to infinity

OpenStudy (anonymous):

so it does not converge

OpenStudy (anonymous):

I get it now. Since cos (npi/2) oscillates between -/+ 1 as n tends to infinity. There is no unique value. Therefore limit does not exist and the sequence is divergent. Am I right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you Sir. [Both of you]

OpenStudy (anonymous):

\[\forall n \in \mathbb{N} , -1\le \cos(\frac{n \pi}{2})\le1 \] So the sequence is always contained in the interval [0;1] But, \[\lim_{n \rightarrow +\infty} 1 =1 \] \[\lim_{n \rightarrow +\infty} -1 =-1 \] So by using the squeeze theorem one concludes that the sequence doesn't converge. That would be the rigurous way to answer it.

OpenStudy (anonymous):

What about an = ln(n)/sqrt(n). Does it converge or diverge?

OpenStudy (anonymous):

Use L'Hopital's Rule to simplify the problem.

OpenStudy (anonymous):

for ^^ use the p rule if 1/n^p if p is greater than one it will diverge and if it is less than one it will converge it will look like this 1/n^1/p to answer your original question, i dont know that one but i know sin and cos dont converge

OpenStudy (anonymous):

They don't, but on paper you have to prove it.